determine the area of a circle having a radius of 243 cm

Found 2 solutions by oberobic, Maths68:
Answer by oberobic(2304)

Area = pi * r^2
Area = pi * 243^2
Area = 59049 * pi cm^2
Area = approximately 185507.905 cm^2
.
Done.

Answer by Maths68(1474)

r=243
pi=3.14
Area of Circle = (pi) * r^2
Area of Circle = 3.14 * (243)^2
Area of Circle =(3.14)*(59049)
Area of Circle = 185413.86cm
Area of Circle =1.8541386 km^2
Huge circle !!!